### v1

@@ III-C Baselines
[…]
@@
IV Results
IV-A Comparison to SOTA Methods
IV-A1 CrossTask
TABLE I: Results on Crosstask dataset with time horizon $T=3$ and $T=4$. Numbers in Bold indicate the best results. The arrow ↑ means higher numbers are better.
@@ IV-A1 CrossTask
ActionDiffusion-Multi-Add (Ours) 37.86 65.58 67.54 22.28 59.30 65.80
[…]
TABLE II: Results on Crosstask dataset with time horizon $T\in\{3,4,5,6\}$. Numbers in Bold indicate the best results. The arrow ↑ means higher numbers are better.
@@
ActionDiffusion-Multi-Add (Ours) 37.86 22.28 13.93 9.13
IV-A2 Coin
TABLE III: Results on Coin and NIV datasets with time horizon $T=3$ and $T=4$. Numbers in Bold indicate the best results. The arrow ↑ means higher numbers are better. ActionDiffusion means our method using Multi-Add mask.
@@ IV-A2 Coin
T=3 DDN 13.90 20.19 - 18.41 32.54 -
Ext-GAIL - - - 22.11 42.20 -
[…]
PDPP 21.33 45.62 51.82 30.20 48.45 57.28
ActionDiffusion-Multi (Ours) 23.58 45.24 53.86 32.96 49.01 57.85
T=4 DDN 11.13 17.71 - 15.97 2.73 -
Ext-GAIL - - - 19.91 36.31 -
[…]
PDPP 14.41 44.10 51.39 26.67 46.89 59.45
ActionDiffusion-Multi (Ours) 17.98 44.43 56.00 29.26 47.71 60.90
@@
Table III shows the results on the Coin dataset with time horizon $T=3$ and $T=4$. For $T=3$, we achieved SOTA performance on SR and mSIOU. Although the mAcc is slightly lower than PDPP, our SR is still 2.25% higher than PDPP. This means that ActionDiffusion can learn the temporal dependencies better than PDPP even though the percentage of correctly predicted actions (ignoring the order) is slightly lower. When $T=4$, we have SOTA results on all metrics. Our mAcc is slightly better than PDPP, SR and mSIoU are 3.57% and 4.61% higher. This also shows our method is better at capturing the temporal dependencies.
IV-A3 NIV
@@
The results on the NIV dataset are shown in Table III. We obtain SOTA performance on all metrics in both time horizons. The SR is 2.76% higher than PDPP when $T=3$ and 2.59% higher when $T=4$.
IV-B Ablation Study
IV-B1 Multi-Add Mask VS. Single-Add Mask
@@ IV-B Ablation Study
In Equation (7), we describe the Multi-Add action-aware noise mask. At each time step within the time horizon $T$, we accumulate the embeddings of all previous actions. We want to compare it with Single-Add noise mask described as follows,
$\begin{bmatrix}0&0&...&0\\ g(a_{e_{1}})&g(a_{e_{1}})+g(a_{e_{2}})&...&g(a_{e_{T-1}})+g(a_{e_{T}})\\ 0&0&...&0\end{bmatrix},$ (12)
where at each time step only the previous action embedding is added to the noise mask. Table IV shows the results of the Multi-Add mask and Single-Add mask on all three datasets with time horizon $T=3$ and $T=4$. Multi-Add outperforms Single-Add on all metrics on the Coin dataset and NIV dataset. On the CrossTask dataset, Single-Add only performs slightly better on SR, and Multi-Add is still better on mAcc and mSIoU. Overall, Multi-Add can better capture and preserve the temporal dependencies between actions.
@@
TABLE IV: Comparison between multi-add noise mask and single-add noise mask on Crosstask, Coin and NIV datasets. Numbers in Bold indicate the best results. The arrow ↑ means higher numbers are better.
@@
Crosstask ActionDiffusion-Multi 37.86 65.58 67.54 22.28 59.30 65.80
@@
Coin ActionDiffusion-Multi 23.58 45.24 53.86 17.98 44.43 56.00
@@
NIV ActionDiffusion-Multi 32.96 49.01 57.85 29.26 47.71 60.90
@@
IV-B2 Self-Attention in U-Net
TABLE V: Comparison between the U-Net with self-attention (w attention) and the U-Net without self-attention (w/o attention). Numbers in Bold indicate the best results. The arrow ↑ means higher numbers are better.
@@ IV-B2 Self-Attention in U-Net
Crosstask w attention 37.86 65.58 67.54 22.28 59.30 65.80
w/o attention 37.86 65.54 67.57 22.38 59.18 65.99
Coin w attention 23.58 45.24 53.86 17.98 44.43 56.00
w/o attention 23.14 44.80 54.00 18.06 44.83 56.32
NIV w attention 32.96 49.01 57.85 29.26 47.71 60.90
w/o attention 31.85 46.30 56.37 27.95 46.52 58.55
We study the effect of the self-attention mechanism in the U-Net in this section. We use the Multi-Add noise mask and test the U-Net with and without self-attention on all three datasets. The results are shown in Table V. ActionDiffusion with and without self-attention achieve comparable results on the CrossTask dataset. On the Coin dataset, ActionDiffusion with and without self-attention also have comparable results. ActionDiffusion with self-attention performs better on SR and mACC when $T=3$. On the NIV dataset, ActionDiffusion with self-attention outperforms the one without self-attention on all metrics with both time horizons. Overall, ActionDiffusion with self-attention performs better than without self-attention on the NIV dataset, while they have comparable performance on the other two datasets. We interpret the reason for this as the size of NIV is much smaller than the other two. ActionDiffusion with attention can better learn the temporal dependencies integrated into the noise mask when the data is limited.
V Conclusion

### v2

@@ 4.4 Baselines
  • P3IV [34] uses transformers with a memory block and a generative module trained with adversarial loss was used to sample the action plan.
@@
5 Results
Table 2: Results on CrossTask dataset with time horizon $T=3$ and $T=4$. Numbers in Bold indicate the best results. The arrow ↑ means higher numbers are better.
@@ 5 Results
ActionDiffusion-MultiAdd (Ours) 37.79 65.38 67.45 22.43 59.42 66.04
5.1 Action Embedding in Noise-Adding
Figure 5: The distributions of diffusion model input after $N$ steps of noise-adding for time horizon $T=3$. Each column shows the distributions at $a_{i},i\in T$. The distribution in blue uses action embedding with Gaussian noise for noise-adding stage. The distribution in orange uses Gaussian noise only. The first row shows the distributions from CrossTask dataset. The second row shows the distributions from Coin dataset. The third row shows the distributions from NIV dataset.
To verify if the noised input follows the normal distribution as mentioned in Section 3.7, we plot the noised actions $\epsilon_{a[1:T]}$ in $x_{n}$, i.e. the second row in Equation 11, for time horizon $T=3$ on all datasets. Additionally, we plot the noised actions $\epsilon_{[1:T]}$ which use $\mathcal{N}(0,\mathbf{I})$ as the noise mask. The time step $N$ is 200, 200 and 50 for CrossTask, Coin and NIV respectively. The results are shown in Figure 5. When using $\epsilon$ as the noise mask for the noise-adding, all noised actions approximately follow the same normal distribution for all three datasets. When using $\epsilon_{a}$ as the noise mask, we can observe from the figure that all action distributions in all datasets also follow normal distributions approximately. The mean of the distribution shifts towards the negative direction for CrossTask and Coin, and shifts toward the positive direction for NIV. For actions $a_{2}$ and $a_{3}$, the means shift further away from zero. The reason is that the mean of action embedding distribution is not zero (Figure 4), $a_{3}$ accumulates the action embeddings of $a_{1}$ and $a_{2}$. Additionally, the standard deviations of $\epsilon_{a}$ are increasing, i.e. $\sigma_{a3}^{2}>\sigma_{a2}^{2},\sigma_{a2}^{2}>\sigma_{a1}^{2}$. The reason is that at $a_{3}$ all previous action embeddings are added to the noise mask.
5.2 Comparison to SOTA Methods
5.2.1 CrossTask
Table 2 shows the results on the CrossTask dataset with time horizon $T=3$ and $T=4$. ActionDiffusion achieves SOTA performances for all metrics in both time horizons. Note that DNN, Ext-GAIL, and P3IV only reported mIoU, and only PDPP reported mSIoU. We also evaluate our method with the longer time horizon $T\in\{3,4,5,6\}$. Table 3 shows the SR for all the time horizons. We again have SOTA performances in all time horizons.
Table 3: Results on CrossTask dataset with time horizon $T\in\{3,4,5,6\}$. Numbers in Bold indicate the best results. The arrow ↑ means higher numbers are better.
@@ 5.2.1 CrossTask
ActionDiffusion-MultiAdd (Ours) 37.79 22.43 13.89 9.66
5.2.2 Coin
Table 4: Results on Coin and NIV datasets with time horizon $T=3$ and $T=4$. Numbers in Bold indicate the best results. The arrow ↑ means higher numbers are better. ActionDiffusion means our method using MultiAdd mask.
@@ 5.2.2 Coin
T=3 DDN [3] 13.90 20.19 - 18.41 32.54 -
Ext-GAIL [2] - - - 22.11 42.20 -
P3IV [34] 15.40 21.67 - 24.68 49.01 -
PDPP [29] 21.33 45.62 51.82 30.20 48.45 57.28
ActionDiffusion-MultiAdd (Ours) 24.00 45.42 54.29 32.96 49.26 57.84
T=4 DDN [3] 11.13 17.71 - 15.97 2.73 -
Ext-GAIL [2] - - - 19.91 36.31 -
P3IV [34] 11.32 18.85 - 20.14 28.36 -
PDPP [29] 14.41 44.10 51.39 26.67 46.89 59.45
ActionDiffusion-MultiAdd (Ours) 18.04 44.54 56.23 29.26 48.14 60.71
@@
Table 4 shows the results on the Coin dataset with time horizon $T=3$ and $T=4$. For $T=3$, we achieved SOTA performance on SR and mSIOU. Although the mAcc is slightly lower than PDPP, our SR is still 2.67% higher than PDPP. This means that ActionDiffusion can learn the temporal dependencies better than PDPP even though the percentage of correctly predicted actions (ignoring the order) is slightly lower. When $T=4$, we have SOTA results on all metrics. Our mAcc is slightly better than PDPP, SR and mSIoU are 4.63% and 4.84% higher. This also shows our method is better at capturing the temporal dependencies.
5.2.3 NIV
@@
The results on the NIV dataset are shown in Table 4. We obtain SOTA performance on all metrics in both time horizons. The SR is 2.76% higher than PDPP when $T=3$ and 2.59% higher when $T=4$. The mACC is 1.18% and 1.39% higher when $T=3$ and $T=4$. The mSIoU is also slightly higher when $T=3$ and $T=4$.
5.3 Ablation Study
5.3.1 Noise Mask
@@
In Equation 7, we describe the MultiAdd action-aware noise mask. At each time step within the time horizon $T$, we accumulate the embeddings of all previous actions. We want to compare it with SingleAdd noise mask described as follows,
$\begin{bmatrix}0&0&...&0\\ g(a_{e_{1}})&g(a_{e_{2}})&...&g(a_{e_{T}})\\ 0&0&...&0\end{bmatrix},$ (12)
where at each time step only one action embedding is added to the noise mask.
Table 5 shows the results of the MultiAdd mask, SingleAdd mask and without mask (NoMask) on all three datasets with time horizon $T=3$ and $T=4$. MultiAdd outperforms SingleAdd and NoMask on all metrics on the Coin dataset and NIV dataset for both time horizons. On the CrossTask dataset, SingleAdd performs the best on SR when $T=3$. NoMask performs the best on mAcc and MSIoU for $T=3$ and all metrics for $T=4$. Although NoMask has the best overall performance on CrossTask, the results of MultiAdd, SingleAdd and NoMask are comparable. Additionally, the performances of NoMask are worse on Coin and NIV, especially on Coin. We interpret the reason for the performance differences is that the differences in action label compositions in action sequences from datasets. Figure 6 shows the distributions of ground truth action labels in action sequences. The action labels in action sequences in CrossTask are more scattered than in Coin and NIV. For instance, lots of action sequences look like $a_{1:T}=[56,0,57]$, where 56, 0 and 57 are the action labels. And most of the action sequences in Coin and NIV look like $a_{1:T}=[48,49,50]$. Overall the distribution of action sequences is more linear in Coin and NIV. We think this is the reason that the action-aware mask works better since it is easier for the action-aware mask to build temporal dependencies. Overall, MultiAdd can perform better when the distribution of action labels is more linear (Coin and NIV). For the more scattered distributions, MultiAdd and SingleAdd are not as effective. Nevertheless, MultiAdd still achieves the SOTA performances.
Figure 6: Distributions of action labels for $T=3$. Each point in the 3D coordinate represents one action sequence. The coordinates in x, y and z directions is the action labels for $a_{1}$, $a_{2}$ and $a_{3}$ respectively.
@@
Table 5: Comparison between MultiAdd noise mask, SingleAdd noise mask and without mask on Crosstask, Coin and NIV datasets. Numbers in Bold indicate the best results. The arrow ↑ means higher numbers are better.
@@
CrossTask ActionDiffusion-Multi 37.79 65.38 67.45 22.43 59.42 66.04
@@
ActionDiffusion-NoMask 37.92 65.53 67.65 22.99 59.48 66.23
Coin ActionDiffusion-Multi 24.00 45.42 54.29 18.04 44.54 56.23
@@
ActionDiffusion-NoMask 12.46 36.08 43.44 2.47 26.96 34.70
NIV ActionDiffusion-Multi 32.96 49.26 57.84 29.26 48.14 60.71
@@
ActionDiffusion-NoMask 29.26 44.81 54.49 21.40 35.59 51.38
5.3.2 Self-Attention in U-Net
Table 6: Comparison between the U-Net with self-attention (w attention) and the U-Net without self-attention (w/o attention). Numbers in Bold indicate the best results. The arrow ↑ means higher numbers are better.
@@ 5.3.2 Self-Attention in U-Net
CrossTask w attention 37.79 65.38 67.45 22.43 59.42 66.04
w/o attention 37.75 65.47 67.45 22.56 59.17 66.15
Coin w attention 24.00 45.42 54.29 18.04 44.54 56.23
w/o attention 22.88 44.52 53.56 18.36 44.88 56.49
NIV w attention 32.96 49.26 57.84 29.26 48.14 60.71
w/o attention 32.22 48.52 57.58 28.82 46.07 59.12
We study the effect of the self-attention mechanism in the U-Net in this section. We use the Multi-Add noise mask and test the U-Net with and without self-attention on all three datasets. The results are shown in Table 6. ActionDiffusion with and without self-attention achieve comparable results on the CrossTask dataset. On the Coin dataset, ActionDiffusion with self-attention performs better on all metrics when $T=3$. ActionDiffusion without self-attention performs better on all metrics when $T=4$, although the results of ActionDiffusion with and without self-attention are comparable. On the NIV dataset, ActionDiffusion with self-attention outperforms the one without self-attention on all metrics with both time horizons. Overall, ActionDiffusion with self-attention performs better than without self-attention on the NIV dataset, while they have comparable performance on the other two datasets. We interpret the reason for this as the size of NIV is much smaller than the other two. ActionDiffusion with attention can better learn the temporal dependencies integrated into the noise mask when the data is limited.
6 Conclusion